# Predicting Case Suffixes With Activity Start and End Times: A Sweep-Line Based Approach

### v1

@@ 4 Evaluation
BPI17W 30,270 240,854 8 7.96 66.18% 12.66 days 286.07 days 706.70%
BPI12W 8,616 59,302 6 6.88 104.11% 8.91 days 85.87 days 369.44%
INS 1,182 23,141 9 19.58 74.27% 70.93 days 599.9 days 459.36%
ACR 954 4,962 16 5.2 32.13% 14.89 days 135.84 days 157.25%
MP 225 4,503 24 20.01 93.78% 20.63 days 87.5 days 130.40%
CVS 10,000 103,906 15 10.39 11.00% 7.58 days 21.0 days 642.64%
CFS 1,000 21,221 29 26.53 54.60% 0.83 days 4.09 days 83.52%
CFM 2,000 44,373 29 26.57 55.41% 0.76 days 5.83 days 83.82%
P2P 608 9,119 21 15 54.43% 21.46 days 108.31 days 78.04%
Real-Life and Synthetic Datasets:
We evaluate our approach using nine event logs [^2] All logs, including a mix of large- and small-scale processes, are available at supplementary material https://shorturl.at/kJA01, selected for their diverse control-flow structures and temporal characteristics. Table 3 reports for each log the number of cases(#Cases), activity instances (#Act-Ins), and unique activities (#Act). (#Act/Case) is the average number of activities per case. CV Len and CV Dur denotes the coefficient of variation in case length and duration respectively [^3] The coefficient of variation (CV) is computed as: $\text{CV}=\frac{\sigma}{\mu}$, where $\sigma$ is the standard deviation and $\mu$ is the mean of case length or duration.. Avg.Dur and Max.Dur shows average and maximum case durations in days.
@@
All logs include start and end timestamps, as required by our approach. The BPI12W and BPI17W logs capture real-life financial processes from a Dutch institution, with BPI17W being a refined version of BPI12W. [^4] https://doi.org/10.4121/uuid:3926db30-f712-4394-aebc-75976070e91f, https://doi.org/10.4121/uuid:5f3067df-f10b-45da-b98b-86ae4c7a310b These logs represent human-performed activities with moderate process complexity. The INS log represents an insurance claims process and features long case durations and high variability. The ACR log, sourced from a Colombian university’s BPM system, models academic credential recognition, with medium trace lengths and a balanced activity structure. The MP (Manufacturing Production) log captures production operations from an ERP system, characterized by long traces and high activity diversity.
@@
We also include three synthetic logs that simulate real-life operational settings. The CVS log models a retail pharmacy process, based on the simulation in Fundamentals of Business Process Management [5], and serves as a large-scale training set. The CFS and CFM logs are anonymized datasets derived from a confidential process, representing small- and medium-scale versions of the same underlying workflow, both featuring high activity density and resource contention. Finally, the P2P (Purchase-to-Pay) log is a synthetic dataset of a procurement process, offering high structural complexity and longer case durations. Table 3 summarizes the statistics of all logs used in the evaluation.
Measures of Goodness:
Baselines:
@@
The experiments compare the performance of our MM architecture against several variants of a Single Model architecture (SM) baselines adapted from Camargo et al. [4]. Specifically, we evaluate four SM variants: Shared Categorical and Full Shared, each tested with and without inter-case features. In the Shared Categorical variant, only the embeddings of categorical inputs (e.g., activity labels) are shared across tasks, while the remaining layers are task-specific. In contrast, the Full Shared variant employs a fully shared architecture, where both categorical and continuous inputs are processed jointly through the same network layers for all tasks. These configurations allow us to isolate the effects of architectural sharing and context-aware features on predictive performance.
@@
To ensure a fair comparison, all models use the same feature encoding and sequence construction methodology described in (Sec. 3.1). Intra-case features are computed individually per case, while inter-case features are derived by sequentially traversing all cases to capture dynamic workload conditions. Activities are encoded as categorical variables using low-dimensional embeddings, and input sequences are constructed via an n-gram strategy, ensuring standardized and consistent feature representation across models.
Experimental Setup:
In SM-based experiments, activity instances that start and finish before the cutoff are excluded from the test set, as SM predicts the next activity, inter-start time, and processing time together but cannot handle ongoing instances. In contrast, MM includes ongoing instances since it can predict their processing time given its activity and start timestamp.
Experiments were conducted on a desktop with an NVIDIA RTX 3090 GPU, Intel Core i9 CPU, and 64 GB RAM. Training the three MM models ($\alpha$, $\beta$, $\gamma$) takes 2.5 hours for smaller logs (e.g., ACR, CFS) and 5-6 hours for larger ones (e.g., CVS, BPI2017W). During testing, the sweep-line approach predicts suffixes of ongoing cases at runtime. For BPI2017W, the approach generated 15,000 predicted events in 100 seconds at 150 events per second.
Next Activity Prediction in the MM Architecture:
A core component of the MM architecture is a model $\alpha$ that predicts the next activity in a case. Using an LSTM-based approach, the model generates multiple possible next activities, each assigned a probability score. A sampling method then selects the next activity. We evaluated MM’s accuracy with three sampling methods: Argmax, which selects the most probable activity [19]; Random Choice, which randomly selects an activity based on the probability distribution [4]; and Daemon Action, a heuristic-based method [1]. All three methods performed similarly, but Daemon Action produced the most consistent results across datasets. Therefore, we selected Daemon Action as the preferred sampling method for model $\alpha$.
Table 4: Hyperparameter Optimization
Parameter - LSTM Explanation Search Space
Batch size # of samples to be propagated [32, 64, 128]
Normalization method Preprocessing - Scaling [lognorm, max]
Epochs # of training epochs 200
N_size Size of the n-gram [5, 10, 15, 20, 25]
L_size LSTM layer sizes [50, 100, 150]
Activation Activation function (hidden layers) [selu, tanh]
Optimizer Weight optimizer [Nadam, Adam, SGD, Adagrad]
Hyperparameter Optimization:
We applied hyperparameter optimization within the sets of values in Table 4 for selecting the most suitable three MM models ($\alpha$ for next activity, $\beta$ for inter-start time, $\gamma$ for processing time). Batch size is optimized for computational efficiency and gradient stability, while normalization methods like log normalization and max scaling improve convergence. The n-gram size (N_size) ensures meaningful historical dependencies, and different LSTM layer sizes (L_size) balance model complexity and training efficiency. Activation functions (Selu, Tanh) help maintain stable gradients, and optimizers (Nadam, Adam, SGD, Adagrad) are chosen for optimal convergence. The training set is split (80% training, 20% validation), and MM is trained with diverse hyperparameter combinations. We optimized hyperparameters using a random search over 50 iterations, which provides a practical balance between search space coverage and computational efficiency. Each model was trained for up to 200 epochs with early stopping(patience = 10) to prevent overfitting and reduce unnecessary training time. These settings are consistent with prior studies [4, 1] and were empirically validated to ensure stable convergence. During optimization, we evaluated performance using two metrics: Mean Absolute Error (MAE)  [^5] MAE is defined as $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ for $\delta^{\text{start}}_{c,k}$ and $\delta^{\text{proc}}_{c,k}$ predictions; and Categorical Cross-Entropy Loss [^6] Categorical Cross-Entropy Loss is defined as $\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{n}\sum_{j=1}^{C}y_{ij}\log(\hat{y}_{ij})$ for next-activity prediction.
To evaluate BiLSTM configuration impact, we tested multiple MM variants with varying BiLSTM layer sizes and n-gram window sizes (see Table 4). Larger models marginally improved accuracy but increased training time and reduced throughput. The default configuration (100-unit layer size, 10 n-gram size, 64 batch size, 200 epochs, Adam optimizer) strikes a practical balance between accuracy and efficiency, offering an effective trade-off for predicting case suffixes in business processes.
Results and Interpretation:
@@
Regarding EQ1, we observe in Fig. 3 that MM outperforms SM-based baselines on the control-flow metric. In the heatmap, we see that the outperformance of MM is most visible for event logs with fewer distinct activities (BPI2017W, BPI2012W, CVS, INS, and ACR, which have less than 20 activities). When the number of activities is lower, the embeddings capture richer information about the sequential relations between activity pairs. The next-activity sampling method used in MM is then able to better exploit this information relative to the argmax next-activity sampling method used in the baselines.
@@
Regarding EQ2, Fig. 4 box plot shows that the three techniques using inter-case features achieve slightly lower inter-case time MAE than the two baselines that rely only on intra-case features. The heatmap in Fig. 4b confirms this observation, particularly for event logs with fewer unique activities (BPI2017W, CVS, INS, and ACR). [^7] An exception is BPI2012W, where all the techniques exhibit similar performance. For logs with larger number of distinct activities, we cannot make a clear conclusion. In the P2P log, the two baselines without inter-case features perform poorly, but in the CFS log, the use of inter-case features does not lead to lower inter-start time MAE and, in fact, MM has the worst performance among all techniques.
Regarding EQ3, Fig. 5 shows that inter-case features have minimal impact on processing time prediction. This is expected since processing time is determined by the activity’s complexity rather than workload, and the inter-case features capture workload information. The heatmap in Fig. 5b confirms this observation. It shows that MM outperforms the baselines in four of the five logs that have less than 20 distinct activities (BPI2012W, CVS, INS, and ACR), but there is no clear pattern in the remaining logs. The outperformance of MM on these four logs is attributable to the fact that it is able to better predict the next-activity, which in turn results in better predictions of the next activity’s processing time.
@@
Additionally, logs like INS and MP contain longer and more variable traces (mean case length $\approx$ 20 activities, CV of case length >> 65%, with INS having durations from 1 to nearly 600 days and a duration CV of 459%), favoring MM’s modular design in learning decoupled control-flow and timing patterns. MM also excels on BPI2017W despite a shorter mean case length ($\approx$ 7 activities), due to its extremely high duration variability (CV of duration 706%). In contrast, CFS and CFM exhibit short (CV of case length << 60%) and homogeneous traces (mean duration << 1 day, CV of duration 83%), where low variability and limited concurrency make simpler SM architectures competitive.
@@
To quantify the benefit of our proposed approach, we compared MM against the best-performing SM baseline for each metric. MM achieves a 6.89% reduction in DL distance, indicating improved control-flow prediction accuracy. For inter-processing time prediction, MM reduces the mean absolute error (MAE) by 24.44%, and for inter-start time prediction, by 18.35%. These results confirm that MM not only improves predictive performance across all dimensions but does so with substantial error reductions over the most competitive baseline.

### v2

@@ 4.1.4 Datasets
BPI17W 30,270 240,854 8 7.96 66.18% 12.66 286.07 706.70%
BPI12W 8,616 59,302 6 6.88 104.11% 8.91 85.87 369.44%
INS 1,182 23,141 9 19.58 74.27% 70.93 599.9 459.36%
ACR 954 4,962 16 5.2 32.13% 14.89 135.84 157.25%
MP 225 4,503 24 20.01 93.78% 20.63 87.5 130.40%
CVS 10,000 103,906 15 10.39 11.00% 7.58 21.0 642.64%
CFS 1,000 21,221 29 26.53 54.60% 0.83 4.09 83.52%
CFM 2,000 44,373 29 26.57 55.41% 0.76 5.83 83.82%
P2P 608 9,119 21 15 54.43% 21.46 108.31 78.04%
GOV 46300 259800 142 5.61 82.00% 143.26 1515.78 124.00%
WorkOrder 19800 149600 24 7.56 48.00% 9.1 112.23 149.00%
P2PFin 4000 30000 16 7.58 67.00% 58.09 697.62 151.00%
@@
All logs include start and end timestamps, as required by our approach. The BPI12W and BPI17W logs capture real-life financial processes from a Dutch institution, with BPI17W being a refined version of BPI12W. [^10] https://doi.org/10.4121/uuid:3926db30-f712-4394-aebc-75976070e91f, https://doi.org/10.4121/uuid:5f3067df-f10b-45da-b98b-86ae4c7a310b These logs represent human-performed activities with moderate process complexity. The INS log represents an insurance claims process and features long case durations and high variability. The ACR log, sourced from a Colombian university’s BPM system, models academic credential recognition, with medium trace lengths and a balanced activity structure. The MP (Manufacturing Production) log captures production operations from an ERP system, characterized by long traces and high activity diversity. The GOV event log records the execution of an approval application system of a government agency. The Workorder is an event log of a field services process at a utilities company. Field services is a type of process in which services are provided to a customer in response to a complaint registered by that customer.
@@
We also include three synthetic logs that simulate real-life operational settings. The CVS log models a retail pharmacy process, based on the simulation in Fundamentals of Business Process Management [9], and serves as a large-scale training set. The CFS and CFM logs are anonymized datasets derived from a confidential process, representing small- and medium-scale versions of the same underlying workflow, both featuring high activity density and resource contention. Finally, the P2P (Purchase-to-Pay) log is a synthetic dataset of a procurement process, offering high structural complexity and longer case durations. Similarly, P2PFin is another anonymized payment process used at a financial institution. Table 4 summarizes the statistics of all logs used in the evaluation.
4.2 Comparison of $M$ and SM Architectures
4.2.1 Baselines
@@
For EQ1, the experiments compare the performance of our $M$ architecture against several variants of a Single Model architecture (SM) baselines adapted from Camargo et al. [5]. Specifically, we evaluate four SM variants: Shared Categorical and Full Shared, each tested with and without resource contention features. In the Shared Categorical variant, only the embeddings of categorical inputs (e.g., activity labels) are shared across tasks, while the remaining layers are task-specific. In contrast, the Full Shared variant employs a fully shared architecture, where both categorical and continuous inputs are processed jointly through the same network layers for all tasks. These configurations allow us to isolate the effects of architectural sharing and context-aware features on predictive performance.
@@
To ensure a fair comparison, all models use the same feature encoding and sequence construction methodology described in (Sec. 3.1). Intra-case features are computed individually per case, while resource contention features are derived by sequentially traversing all cases to capture dynamic workload conditions. Activities are encoded as categorical variables using low-dimensional embeddings, and input sequences are constructed via an n-gram strategy, ensuring standardized and consistent feature representation across models.
4.2.2 Results and Interpretation
@@
Regarding EQ1(a), we observe in Fig. 3 that $M$ outperforms SM-based baselines on the control-flow metric. In the heatmap, we see that the outperformance of $M$ is most visible for event logs with fewer distinct activities (BPI2017W, BPI2012W, CVS, INS, and ACR, which have less than 20 activities). When the number of activities is lower, the embeddings capture richer information about the sequential relations between activity pairs. The next-activity sampling method used in $M$ is then able to better exploit this information relative to the argmax next-activity sampling method used in the baselines.
@@
Regarding EQ1(b), Fig. 4 box plot shows that the three techniques using resource contention features achieve slightly lower inter-start time MAE than the two baselines that rely only on intra-case features. The heatmap in Fig. 4b confirms this observation, particularly for event logs with fewer unique activities (BPI2017W, CVS, INS, and ACR). [^11] An exception is BPI2012W, where all the techniques exhibit similar performance. For logs with larger number of distinct activities, we cannot make a clear conclusion. In the P2P log, the two baselines without resource contention features perform poorly, but in the CFS log, the use of resource contention features does not lead to lower inter-start time MAE and, in fact, $M$ has the worst performance among all techniques.
Regarding EQ1(c), Fig. 5 shows that resource contention features have minimal impact on processing time prediction. This is expected since processing time is determined by the activity’s complexity rather than workload, and the resource contention features capture workload information. The heatmap in Fig. 5b confirms this observation. It shows that $M$ outperforms the baselines in four of the five logs that have less than 20 distinct activities (BPI2012W, CVS, INS, and ACR), but there is no clear pattern in the remaining logs. The outperformance of $M$ on these four logs is attributable to the fact that it is able to better predict the next-activity, which in turn results in better predictions of the next activity’s processing time.
@@
Additionally, logs like INS and MP contain longer and more variable traces (mean case length $\approx$ 20 activities, CV of case length >> 65%, with INS having durations from 1 to nearly 600 days and a duration CV of 459%), favoring $M$’s modular design in learning decoupled control-flow and timing patterns. $M$ also excels on BPI2017W despite a shorter mean case length ($\approx$ 7 activities), due to its extremely high duration variability (CV of duration 706%). In contrast, CFS and CFM exhibit short (CV of case length << 60%) and homogeneous traces (mean duration << 1 day, CV of duration 83%), where low variability and limited concurrency make simpler SM architectures competitive.
@@
To quantify the benefit of our proposed approach, we compared $M$ against the best-performing SM baseline for each metric. $M$ achieves a 6.89% reduction in DL distance, indicating improved control-flow prediction accuracy. For inter-processing time prediction, $M$ reduces the mean absolute error (MAE) by 24.44%, and for inter-start time prediction, by 18.35%. These results confirm that $M$ not only improves predictive performance across all dimensions but does so with substantial error reductions over the most competitive baseline.
4.3 Impact of Resource Availability Features
We designed two experiments to evaluate different aspects of EQ2. The first experiment (EQ2.1) investigates whether the proposed resource availability features improve the performance of the $M$ approach using a synthetic event log where resources follow predefined (fixed) working calendars. The second experiment (EQ2.2) examines the impact of incorporating resource availability features on the predictive performance of the $M$ approach when applied to real-life event logs.
4.3.1 Baselines
For the evaluation of EQ2, we used the base $M$ approach without resource availability features as the baseline. This baseline represents the standard configuration of our approach, relying solely on inter and intra-case features without considering resource availability $R_{avail}$. By comparing this baseline with the extended $M$ variant $MM_{avail}$ that incorporates resource availability features, we can directly assess the contribution of these features to predictive performance. This design allows us to isolate and quantify the effect of resource-related information on control flow, inter-start time, and processing time predictions across both synthetic and real-life event logs.
4.3.2 Datasets
To evaluate EQ2.1, we employed a simulation model of a loan application process as the baseline. From this model, we generated eight synthetic event logs, each corresponding to a different resource-availability scenario, as outlined in Table 5.
In the first scenario (a), all resources are continuously available, operating 24 hours a day, seven days a week. This results in a non-intermittent, dense, and homogeneous workload. In the second scenario (b), half of the resources share the same non-intermittent, dense schedule, working 24 hours a day from Monday to Sunday. The remaining half of the resources work 22 hours a day, Monday to Sunday, creating a heterogeneous workload. In the third scenario (c), all resources work 16 hours a day, Monday to Sunday, producing an intermittent, dense, and homogeneous workload. To create a heterogeneous workload in the fourth scenario (d), half of the resources work 16 hours a day, while the other half work 12 hours a day, Monday to Friday, resulting in an intermittent, dense, and heterogeneous calendar. To generate a sparse calendar, the resource workload should be less than 50 hours per week. In scenario (e), a resource works 8 hours a day, Monday to Friday, resulting in an intermittent, sparse, and homogeneous workload. In scenario (f), the workload is made intermittent by having half of the resources work 8 hours a day, while the other half work 4 hours a day, Monday to Friday. This produces an intermittent, sparse, and heterogeneous calendar. For a non-intermittent, sparse, and homogeneous workload, scenario (g) features a calendar where all resources work 8 hours a day on Mondays and Fridays only. In scenario (h), half of the resources work 6 hours a day on Tuesdays, Thursdays, and Saturdays, while the other half work 6 hours a day on Mondays, Wednesdays, and Fridays, creating an intermittent, sparse, and heterogeneous calendar. For each of these eight scenarios, we used the Apromore simulator to generate a synthetic event log consisting of 2,000 cases. The inter-arrival times between cases followed a Poisson distribution with a mean inter-arrival time of 20 minutes. Arrivals only occurred within the working hours defined by the corresponding resource calendars.
To evaluate EQ2.2, we used the same real-life event logs as in EQ1. The $M$ baseline without resource availability features was compared against the extended $M$ variant that incorporates these features $MM_{avail}$ to assess the impact on control flow, inter-start time, and processing time predictions.
Table 5: Resource Availability Scenarios
Scenario Resource Availability Workload Type
a 24 hours a day, 7 days a week Non-intermittent, dense, homogeneous
b Half resources: 24 hours a day, Monday to Sunday; half resources: 22 hours a day, Monday to Sunday Non-intermittent, dense, heterogeneous
c 16 hours a day, Monday to Sunday Intermittent, dense, homogeneous
d Half resources: 16 hours a day, Monday to Friday; half resources: 12 hours a day, Monday to Friday Intermittent, dense, heterogeneous
e 8 hours a day, Monday to Friday Intermittent, sparse, homogeneous
f Half resources: 8 hours a day, Monday to Friday; half resources: 4 hours a day, Monday to Friday Intermittent, sparse, heterogeneous
g 8 hours a day, Monday and Friday each week Non-intermittent, sparse, homogeneous
h Half resources work 6 hours a day on Tuesday, Thursday, and Saturday; half resources work 6 hours a day on Monday, Wednesday, and Friday Non-intermittent, sparse, heterogeneous
4.3.3 Results and Interpretation
Regarding EQ2.1, Table 6 presents the results for goodness measures (DL Distance, MAE Inter-start Time, and MAE Processing Time) for each resource availability scenario presented in Table 5 for modified loan application process. These results provide key insights into how varying work shift scenarios for available resources affect the control-flow and temporal accuracy of case suffix prediction for $M$ and $MM_{avail}$.
The best performance in terms of DL Distance and MAE Inter-start Time is observed in Scenario (f), where resources are intermittently available, with half working 8 hours a day, Monday to Friday, and the other half working 4 hours a day, Monday to Friday. This scenario achieves the lowest values for DL Distance (0.33) and MAE Inter-start Time (3.81). Additionally, MAE Processing Time is also the best in this scenario (1.64). These results suggest that intermittent, sparse, and heterogeneous resource availability improves prediction accuracy. The model benefits from clear on-off-duty cycles, reducing congestion and enabling more efficient task scheduling.
Conversely, Scenario (a), where resources are continuously available 24/7, yields higher DL Distance (0.43) and MAE Inter-start Time (3.97), along with a relatively high MAE Processing Time (1.73). This occurs because continuous availability leads to resource congestion and task overlap, which reduces the efficiency of load distribution and task management. The absence of structured work–rest cycles limits the model’s ability to learn meaningful patterns, and as a result, resource availability features have zero impact in this scenario because they do not vary over time and therefore provide no informative signal to the predictive model.
Scenarios with fixed working hours, like Scenario (c) with 16 hours per day, show intermediate performance. Specifically, Scenario (c) results in DL Distance (0.39), MAE Inter-start Time (3.87), and MAE Processing Time (1.68). While this schedule improves model performance compared to continuous availability, it still leads to higher errors compared to the more heterogeneous and sparse schedules, such as Scenario (f).
Scenario (e), with a sparser schedule of 8 hours per day, Monday to Friday, results in DL Distance (0.37) and MAE Inter-start Time (3.85), while maintaining lower processing times (1.68) compared to other scenarios. This indicates that a predictable, sparse resource schedule improves prediction accuracy, as the model can better anticipate resource availability, leading to more optimal scheduling.
The results indicate that intermittent and sparse resource schedules, particularly those with work hours distributed across blocks, improve prediction accuracy by 15.38% for DL Distance, 3.05% for inter-start time, and 3.53% for processing time. Incorporating resource availability features therefore leads to measurable gains in the accuracy of predicting the next activity, inter-start times, and processing times.
Table 6: EQ.2.1 - loan application process results
Scenerio DL Distance MAE Inter-start Time MAE Processing Time
$M$ $_{avail} MM MM$$_{avail} MM$ $MM_{avail}$
a 0.43 0.43 3.97 3.97 1.73 1.73
b 0.41 0.40 3.94 3.92 1.73 1.72
c 0.43 0.39 3.91 3.87 1.71 1.68
d 0.40 0.37 3.90 3.84 1.70 1.66
e 0.41 0.37 3.92 3.85 1.72 1.68
f 0.39 0.33 3.93 3.81 1.70 1.64
g 0.42 0.34 3.92 3.84 1.71 1.66
h 0.43 0.37 3.93 3.86 1.72 1.67
Regarding EQ2.2, Table 7 presents a comparison between $M$ and $MM_{avail}$ for predicting case suffixes in an ongoing business process. The empirical results demonstrate that $MM_{avail}$ consistently outperforms the baseline across all three evaluation metrics: DL Distance, MAE Inter-start Time, and MAE Processing Time. Notably, $MM_{avail}$ reduces the DL Distance, for instance, from 0.48 to 0.39 in the BPI2012W dataset, representing a substantial improvement of 18.75%. This significant reduction indicates a refined capability to capture control-flow fluctuations. This improvement is especially noticeable in datasets with variable resource schedules, such as INS, MP, and P2P. $MM_{avail}$ is able to account for resource availability gaps by incorporating resource availability features, providing a more accurate prediction. The table also highlights that these real-life logs tend to have less dense(Den.) and more sparse resource calendars, with the highest incidence of intermittent(Int.) schedules. Moreover, the availability features such as Time Until Next On-Duty and Time Since Last Off-Duty enable the $MM_{avail}$ model to better understand resource on-off-duty patterns. Consequently, this mitigates sequence alignment errors, yielding superior chronological precision for the control flow of ongoing cases.
In terms of MAE Inter-start Time, $MM_{avail}$ consistently outperforms $M$, with significant improvements observed in datasets such as INS (from 53.05 to 45.04) and ACR (from 43.98 to 36.02). These datasets feature highly intermittent and sparse resource availability calendars, where accurately predicting the time between task completions requires an understanding of resource on-off-duty cycles. The resource availability features of $MM_{avail}$ enable it to effectively manage these intermittent and sparse schedules, leading to more precise predictions of when the next task should begin. In contrast, $M$ struggles with such schedules, as it does not account for resource availability, resulting in less accurate predictions of inter-start times for the predicted case suffixes.
For the MAE Processing Time, both $M$ and $MM_{avail}$ perform similarly on certain logs, such as BPI2017W and CVS. This similarity arises because processing time is less influenced by workload factors like resource availability. Instead, it is primarily determined by the nature of the activity itself. However, $MM_{avail}$ shows clear advantages on datasets like Work Order, INS, and P2PFin, where the coefficient of variation (CV) for case length varies between 48% and 74%, and the CV for case duration varies between 149% and 459%. On other logs, the improvement remains either consistent or improved with respect to the $M$ approach. Hence, the inclusion of availability features enhances the model’s ability to predict task durations, factoring in the on-off-duty cycles of resources. $MM_{avail}$ demonstrates better performance, particularly on event logs with sparse and highly intermittent resource on-off-duty schedules, where the availability features lead to more accurate predictions compared to the $M$ approach.
In conclusion, $MM_{avail}$ demonstrates superior performance compared to the MMMM approach for all three metrics. This improvement is fundamentally linked to the accuracy of start-time predictions, which governs the correctness of the DL distance. Accurate start times are essential because the model processes activity instances based on their sorted chronological order. For instance, consider two overlapping activities: one starting at 10 and another at 12. If the start time of the second activity is underestimated (e.g., predicted as 9), the sorting logic inevitably inverts their sequence. Such disordering negatively impacts the sweep-line based technique used by the model. The $MM_{avail}$ method addresses this by leveraging resource availability features to refine start-time estimates, thereby preserving the correct sequence of activities and significantly improving both control flow accuracy and processing time prediction.
Table 7: EQ.2.2 - Comparison of $M$ and $MM_{avail}$ Approaches
Dataset Den. Int. DL Distance MAE Inter-start Time MAE Processing Time
$M$ $_{avail} MM MM$$_{avail} MM$ $MM_{avail}$
BPI2017W 33.04% 0.36 0.45 0.43 15.44 14.01 0.04 0.04
BPI2012W 26.36% 0.43 0.48 0.39 13.06 11.05 0.06 0.05
CVS 39.31% 0.27 0.50 0.50 1.52 1.43 0.01 0.01
INS 10.82% 0.36 0.49 0.44 53.05 45.04 2.66 2.43
ACR 3.04% 0.23 0.28 0.26 43.98 36.03 0.02 0.02
MP 28.27% 0.45 0.35 0.27 28.74 24.05 1.16 1.12
CFS 43.41% 0.25 0.53 0.53 0.69 0.63 0.43 0.39
CFM 34.44% 0.24 0.50 0.48 0.60 0.55 0.64 0.61
P2P 32.01% 0.31 0.56 0.48 67.94 64.82 1.39 1.37
GOV 32.96% 0.41 0.34 0.28 33.58 27.98 1.78 1.76
Work Order 21.13% 0.48 0.53 0.51 12.62 9.76 1.62 1.58
P2PFin 11.31% 0.29 0.28 0.27 21.72 18.43 0.72 0.67